Roll Corporation has 20,000 shares of common stock outstanding. It's financed entirely with equity. The un-levered cost of capital is 14%. The company distributes all of its earnings to equity holders as dividends at the end of each year. And Roll Corporation is subject to a corporate tax rate of 30%.

Roll Corporation estimates that its annual EBIT will be as follow: 20% Bust with EBIT 2600, 50% Expected with EBIT 3620, 30% Expansion with EBIT 4900.

The firm expects the earnings will be unchanged in perpetuity.

A. What is the market value of Roll Corporation?

B. Suppose Roll Corporation issues $8000 of perpetual debt and use the proceeds to repurchase common stock. The bond will sell at par with a 6% annual coupon rate. After the sales of the bonds, Roll Corporation will remain the new capital structure indefinitely.

B1: What is the new value of the firm?

B2. What is the new vale of the firm's equity?

B3. What is the required return on Roll Corporation's equity after restructuring?

B4. Will the presence of corporate tax increase or decrease the firm value? Why?
Answer #1

Answer

A) Given that,

Un-levered cost of capital = 14%

Tax rate = 30%

Annual EBIT = 2600 + 3620 + 4900

= 11120

Market value of firm = EBIT / Cost of capital

= 11120 / 14%

= 55600

B) $8000 debt is issued and used to repurchase stock.

B1) Market value of levered firm,

= Market value of unlevered firm + Tax advantage of debt - equity bought back

= 55600 + 2400 - 8000

= 50000

Tax advantage of debt = Tax rate * debt

= 0.3 * 8000

= 2400

B2) New value of equity = Market value - debt

= 50000 - 8000

= 42000

B3) Required return on equity after restructuring

= PAT / Market value of equity

= 7448 / 42000

= 17.73%

PAT = ( EBIT - Interest ) ( 1- tax)

= [ 11120 - (6% of 8000) ] ( 1- 0.3 )

= 7448

B4) Presence of corporate tax will increase the value of the firm as when debt is employed and

interest is paid on it then the interest is treated as an expense on which tax is saved.